OpenStudy (anonymous):

Slove by Subsitution. .5x +.25y=36 AND y+18=16x

OpenStudy (anonymous):

is that 0.5x + 0.25y=36 or just 5x + 25y=36

OpenStudy (anonymous):

0.5+0.25y=36